A refrigerator requires 20 J of work and exhausts 70 J of heat per cycle. What is the refrigerator's coefficient of performance?
Answer #1

Efficiency, n = 20/70 = 0.286

Also,

Efficiency, n = (Qh - Qc)/Qh

0.286 = (70 - Qc)/70

Qc = 50 J

COP = Qc/W = 50/20 = 2.5

Comment in case any doubt please rate my answer....